OpenStudy (anonymous):

Solve for x: (4overX)+(4overX^2-9)=(3overX-3)

OpenStudy (anonymous):

\[{4 \over x}+{4 \over x^2-9}-{3 \over x-3}={4(x^2-9)+4x-3x(x+3) \over x(x^2-9)}.\] The expression is 0 when the numerator equals 0. You need now to solve the quadratic equation \(4(x^2-9)+4x-3x(x+3)=x^2-5x-36=0\).

OpenStudy (anonymous):

answer?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

x=4 and x=-9?

OpenStudy (anonymous):

x=-4 or x=9.

OpenStudy (anonymous):

thanks